# ToDRE: Visual Token Pruning via Diversity and Task Awareness for Efficient Large Vision-Language Models

### v1

@@ 5.3 Ablation Study
We conduct ablation studies to evaluate individual and combined contributions of two stages in our framework. As shown in Table 4, applying Stage 2 only, which removes all visual tokens at a selected LLM layer without early-stage diversity-aware selection, already reduces the overall inference time by 6.5% compared to unpruned LLaVA-NeXT-7B baseline (from 77:04 to 72:46), while maintaining a strong average performance of 98.9%. The limited efficiency gain is expected, as Stage 2 only accelerates the latter part of inference, and most tasks involve generating very short outputs.
In contrast, applying Stage 1 only pruning, which retains 25% or 10% of tokens based on token diversity, yields substantial time savings of 42% (44:40) and 60% (30:38), respectively, with minimal drops in performance. When incorporating both stages (Stage 1 + Stage 2), we observe consistent improvements: First, at the 25% ratio, performance improves from 97.4% to 97.7% with total time reduced (from 30:38 to 29:51). Second, at the 10% ratio, performance increases from 95.1% to 95.5%, with total time reduced (from 30:38 to 29:31). Overall, ToDRE reduces inference time by 44% and 61.7% at the 25% and 10% retain ratios, respectively, while even improving performance (up to +0.4% gain). These results confirm that the second stage—full visual token removal based on visual-task relevance—provides complementary benefits to the diversity-based Stage 1, leading to improved accuracy-efficiency trade-offs under various compression settings.
@@
FastV [10] 8.2 (↓78.0%) 14.1 (↓11.3%) 2.1 (1.4×) 88.8%
SparseVLM [62] 6.9 (↓78.0%) 14.1 (↓11.3%) 2.5 (1.7×) 86.3%
FasterVLM [61] 6.1 (↓80.6%) 13.6 (↓14.5%) 2.7 (1.8×) 91.4%
GlobalCom2 [39] 6.1 (↓80.6%) 13.9 (↓12.6%) 2.7 (1.8×) 92.2%
ToDRE (Ours) 6.0 (↓80.9%) 13.6 (↓14.5%) 2.8 (1.9×) 95.1%
@@
LLaVA-NeXT-7B [37] 77:04 1519.6 72.0 64.2 87.7 100.0%
Stage 2 only 72:46 1497.5 70.5 63.7 87.7 98.9%
@@
Stage 1 only 44:40 1467.0 70.0 61.6 87.5 97.4%
Stage 1 + Stage 2 (ToDRE) 43:09 1483.3 70.2 61.4 87.7 97.7%
@@
Stage 1 only 30:38 1400.0 69.6 60.1 86.1 95.1%
Stage 1 + Stage 2 (ToDRE) 29:31 1413.1 70.0 60.3 85.9 95.5%
Table 4: Ablation study on two-stage token compression. We evaluated the individual and combined effects of proposed two-stage pruning pipeline under retain ratios of 25% and 10%.
@@ 6 Conclusion
In this work, we systematically analyze redundancy in LVLM inference and identify two key inefficiencies: (1) redundant visual tokens that inflate intra-modal computation, and (2) tokens that contribute little cross-modal information during decoding. To address these, we propose ToDRE, a training-free, architecture-agnostic framework that first selects a maximally diverse subset of visual tokens via greedy $k$-center sampling, then removes the rest once cross-modal attention fades. Experiments on twelve image- and video-language benchmarks show that ToDRE prunes up to 90% of visual tokens while preserving 95.1% accuracy, achieving 2.6$\times$ faster inference and 14.5% lower memory usage than uncompressed baselines.
Appendix
In this Appendix, we first provide a detailed introduction of the benchmarks used in our experiments in Sec. A. Next, we present a theoretical analysis of our two-stage token compression paradigm in Sec. B. Additional ablation studies are reported in Sec. C, where we specifically examine the effects of the pruning threshold and layer selection in  C.1 and Sec. C.2, respectively. We further investigate the negligible decoding-stage attention to visual tokens in Sec. D. In subsequent Sec.E.1, we demonstrate the unique advantages of ToDRE with question-answering cases in a chatbot scenario, targeting those long and free-form responses. Furthermore, in Sec.E.2, we extensively perform qualitative comparisons under seven benchmarks to visualize the differences between our diversity-driven approach and existing attention-driven token pruning method. Finally, we discuss the limitations of this work in Sec.F. Unless otherwise stated, all experiments in this paper were conducted on 4 $\times$ NVIDIA RTX 3090 GPUs.
Algorithm 1 Proposed Greedy $k$-Center Algorithm for Token Retention
Input: $V\in\mathbb{R}^{n\times d}$: visual features; $\alpha\in\mathbb{R}^{n}$ or $\mathbb{R}^{f\times t}$: CLS-to-token attention scores; $k$: number of tokens to retain
@@
5:     $a,b\leftarrow\arg\max_{f,t}\alpha_{f,t}$
6:     $p\leftarrow a\cdot t+b$
7:end if
8:// Stage 2: Greedy k-center Expansion
9:$\mathcal{C}\leftarrow\{p\}$triangleright Initialize with the pivot token
10:$s\leftarrow VV_{p}^{\top}$triangleright Initial similarity: max_sim
11:for $i=1$ to $k-1$ do
12:     $c\leftarrow\arg\min s$triangleright Find most dissimilar token
13:     $\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$
14:     $s\leftarrow\max(s,VV_{c}^{\top})$triangleright Update max similarity
15:end for
16:return $\mathcal{C}$
Appendix A Experimental Details
A.1 Benchmarks
We evaluate our method on a range of widely used benchmarks, collectively designed to assess various aspects of multimodal intelligence. For image understanding tasks, we conduct experiments on eight benchmarks: MME [14], ScienceQA [41], GQA [21], POPE [31], MMBench and MMBench-CN [40], VizWiz [19], and VQAv2 [17]. For video understanding tasks, we evaluate our method on four benchmarks: VideoMME [15], EgoSchema [42], MLVU [64], and LongVideoBench [55].
@@ A.1 Benchmarks
GQA is a benchmark designed to evaluate a model’s structured understanding and reasoning capabilities over visual scenes. It is built upon three key components: images, scene graphs, and carefully constructed questions. Each image is accompanied by a scene graph derived from the Visual Genome dataset [24], which provides detailed object-level annotations, attributes, and relationships within the scene. The questions are automatically generated from these graphs to ensure semantic clarity and logical consistency, enabling fine-grained assessment of a model’s reasoning ability. Following standard practice, we report accuracy on the test-dev set, which contains 12,578 image-question pairs.
@@
POPE is a benchmark designed to assess object hallucination in vision-language models through binary questions about the presence of specific objects in images. The images are sourced from the MSCOCO dataset [34], and evaluation is based on the average F1 score across three sampling strategies, using a total of 8,910 image-question pairs.
@@
MMBench is a hierarchical benchmark designed to comprehensively evaluate multimodal model capabilities across three levels: perception and reasoning (L1), six sub-skills (L2), and 20 specific tasks (L3). Each task consists of multiple-choice questions. The benchmark is available in both English and Chinese, with 4,377 and 4,329 image-question pairs respectively. We use both MMBench and MMBench-CN for evaluation.
@@
VQAv2 is a benchmark designed to evaluate a model’s visual recognition and reasoning capabilities through open-ended questions grounded in real-world images. It contains 265,016 images from the MSCOCO dataset [34], with each image paired with at least three questions. To mitigate bias, the dataset adopts an adversarially balanced design, ensuring that each question appears with multiple images leading to different answers. Each question is annotated with ten human-provided answers. We use the test-dev set for evaluation, which includes 107,394 image-question pairs, with scoring based on standard automatic metrics.
@@
EgoSchema is a diagnostic benchmark designed to evaluate the very long-form video-language understanding capabilities of LVLMs. Derived from the Ego4D dataset [18], it comprises over 5,000 human-curated multiple-choice question-answer pairs spanning more than 250 hours of egocentric video footage, covering a broad range of natural human activities and behaviors. Each question is based on a three-minute-long video clip and requires selecting the correct answer from five options. In our experiments, we evaluate on the EgoSchema test set, which contains 5,031 video-question pairs.
@@ A.2 Backbone Models
LLaVA-NeXT [37] is also referred to as LLaVA-1.6, extending LLaVA-1.5 [36] by introducing a dynamic high-resolution processing strategy that enhances performance in tasks requiring visual reasoning, OCR, and world knowledge. In contrast to the fixed resolution scaling used in LLaVA-1.5, LLaVA-NeXT adaptively adjusts the input resolution by selecting an optimal aspect ratio based on the original image. The resolution can be increased by up to 4×. Notably, this enhancement is achieved without modifying the visual encoder. Instead, each high-resolution image is divided into multiple sub-images of the original size, which are independently encoded and then concatenated before being passed to the language model. All experiments in this study are based on the 7B and 13B versions of LLaVA-NeXT.
A.3 Comparing Methods
@@
FastV [10] is a training-free method that reduces computational overhead in vision-language models by performing early-stage visual token pruning. It identifies and removes the least relevant tokens after the second LLM layer by averaging attention scores.
@@
SparseVLM [62] ranks the importance of both visual and textual tokens based on cross-modal attention, and introduces adaptive sparsity ratios along with a token recycling strategy to better utilize discarded information.
@@
FasterVLM [61] leverages attention from the [CLS] token to visual tokens as an importance indicator, pruning the less relevant visual tokens accordingly.
@@
GlobalCom2 [39] is designed for high-resolution image understanding tasks that receive both a global thumbnail and multiple local crops. The thumbnail provides global contextual guidance to guide the compression of each crop in a task-specific manner.
Appendix B Theoretical Perspective: Orthogonality of Intra- and Cross-Modal Redundancy
@@
$\mathcal{V}=\operatorname{Span}(W_{V}),\qquad\mathcal{T}=\operatorname{Span}(W_{T}),$
@@
$W_{V}^{\top}W_{T}=0\qquad(\mathcal{V}\;\perp\;\mathcal{T}).$
@@
Proof of (11). Set $X_{ij}=\kappa(v_{i},v_{j})$ and $Y_{k}=\rho(v_{k},T)$. By eqs. 8 and 9,
$D_{\kappa}(V)=\frac{1}{N^{2}}\!\sum_{i\neq j}X_{ij},\qquad R_{\rho}(V,T)=\frac{1}{N}\!\sum_{k=1}^{N}Y_{k}.$
@@
Appendix C More Ablation Studies
C.1 Threshold in Relevance-driven Visual Token Reduction
@@
LLaVA-NeXT-7B [37] 77:04 1519.6 72.0 64.2 87.7 100.0%
0.01 73:12 1495.0 70.6 63.7 87.8 98.9%
0.03 (Ours) 72:46 1497.5 70.5 63.7 87.7 98.9%
0.05 72:08 1495.0 68.2 63.7 87.8 98.1%
0.10 60:27 1517.9 65.9 58 87.7 95.4%
Table 5: Ablation study on threshold $\bm{\tau}$ in relevance-driven visual token reduction. When both attention ratios $\alpha_{t\to v}$ and $\alpha_{v\to t}$ (Equation 7) fall below the threshold $\tau$, all remaining visual tokens are removed from the LLM input.
We conduct an ablation study on the threshold $\tau$ used in the relevance-driven visual token reduction strategy, as shown in Table 5. Varying $\tau$ controls the aggressiveness of token pruning based on the measured attention ratios $\alpha_{t\to v}$ and $\alpha_{v\to t}$ (Equation 7). A larger $\tau$ leads to more extreme pruning but may sacrifice accuracy, while a smaller $\tau$ makes token pruning more conservative and thus improves performance at the cost of increased computational burden.
We choose $\tau=0.03$ as our default setting, which yields the optimal trade-off between efficiency and performance—maintaining 98.9% average accuracy while reducing inference time by over 4 minutes compared to the uncompressed baseline.
C.2 Layers in Relevance-driven Visual Token Reduction

### v2

@@ 5.3 Ablation Study
FastV [chen2024fastv] 8.2 (↓73.9%) 14.1 (↓11.3%) 2.1 (1.4×) 88.8%
SparseVLM [zhang2024sparsevlm] 6.9 (↓78.0%) 14.1 (↓11.3%) 2.5 (1.7×) 86.3%
FasterVLM [zhang2024fastervlm] 6.1 (↓80.6%) 13.6 (↓14.5%) 2.7 (1.8×) 91.4%
GlobalCom2 [liu2025globalcom2] 6.1 (↓80.6%) 13.9 (↓12.6%) 2.7 (1.8×) 92.2%
DivPrune [alvar2025divprune] 6.0 (↓80.9%) 13.6 (↓14.5%) 2.8 (1.9×) 93.5%
ToDRE (Ours) 6.0 (↓80.9%) 13.6 (↓14.5%) 2.9 (1.9×) 95.0%
@@
We conduct ablation studies to evaluate individual and combined contributions of the two stages in our framework. As shown in Table˜5, applying Stage 2 only, which removes all visual tokens at a selected LLM layer without early-stage diversity-aware selection, already reduces the overall inference time by 8.8% compared to unpruned LLaVA-NeXT-7B baseline (from 77:04 to 70:15), while maintaining a lossless average performance of 100.0%. The limited efficiency gain is expected, as Stage 2 only accelerates the latter part of inference, and most tasks involve generating very short outputs.
In contrast, applying Stage 1 only, which retains 25% or 10% of tokens based on token diversity, yields substantial time savings of 37.5% (48:10) and 59.4% (31:18), respectively, with minimal drops in performance. When incorporating both stages (Stage 1 + Stage 2), we observe consistent improvements: First, at the 25% ratio, performance improves from 98.8% to 98.9% with total time reduced (from 48:10 to 44:18). Second, at the 10% ratio, performance increases from 95.8% to 96.0%, with total time reduced (from 31:18 to 29:43). Overall, ToDRE reduces inference time by 42.5% and 61.4% at the 25% and 10% token retention ratios, respectively, while even improving performance (up to +0.2% gain). These results confirm that the second stage—full visual token removal based on visual-task relevance—provides complementary benefits to the diversity-based Stage 1, leading to improved accuracy-efficiency trade-offs under various compression settings.
@@
[…]
Stage 2 only 70:15 1522.7 71.9 64.3 87.6 100.0%
@@
Stage 1 only 48:10 1503.8 70.6 63.1 87.5 98.8%
Stage 1 + Stage 2 (ToDRE) 44:18 1504.3 70.7 63.3 87.5 98.9%
@@
Stage 1 only 31:18 1458.6 70.4 59.4 85.0 95.8%
Stage 1 + Stage 2 (ToDRE) 29:43 1469.3 70.5 59.4 85.0 96.0%
Table 5: Ablation study on two‐stage token compression. We evaluated the individual and combined effects of proposed two-stage pruning pipeline under retention ratios of 25% and 10%.
@@ 6 Conclusion
[…]
Supplementary Material
Outline
In this Supplementary Material, we first provide a detailed introduction of the benchmarks used in our experiments in Sec. 1. Next, we present a theoretical analysis of our two-stage token compression paradigm in Sec. 2. Additional ablation studies are presented in Sec. 3, where we further analyze different pivot token selection strategies in Sec. 3.1, as well as the effects of the pruning threshold and layer selection in Sec. 3.2 and Sec. 3.3, respectively. We further investigate the negligible decoding-stage attention to visual tokens in Sec. 4. In subsequent Sec. 5.1, we demonstrate the unique advantages of ToDRE with question-answering cases in a chatbot scenario, targeting those long and free-form responses. Finally, in Sec. 5.2, we perform extensive qualitative comparisons under seven benchmarks to visualize the differences between our diversity-driven approach and existing attention-driven token pruning method. Unless otherwise stated, all experiments in this paper were conducted on 4 NVIDIA RTX 3090 GPUs.
Algorithm 1 Proposed Greedy Max-Sum Diversification for Token Retention
Input: $V\in\mathbb{R}^{n\times d}$: visual features; $\alpha\in\mathbb{R}^{n}$ or $\mathbb{R}^{f\times t}$: [CLS]-to-token attention; $k$: #tokens to retain
@@
5:   $(a,b)\leftarrow\arg\max_{f,t}\alpha_{f,t}$;    $p\leftarrow a\cdot t+b$
6:end if
7:// Stage 2: Greedy Max-Sum Diversification
[…]
12:for $i=1$ to k1k-1 do
13:   $c\leftarrow\arg\min s$ $\triangleright$ pick token with smallest cumulative similarity
14:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$
15:   $s\leftarrow s+XX_{c}^{\top}$ $\triangleright$ update: add similarity to the new token
16:   $s_{c}\leftarrow+\infty$ $\triangleright$ mask the newly selected index
17:end for
18:return $\mathcal{C}$
1 Experimental Details
1.1 Benchmarks
We evaluate our method on a range of widely used benchmarks, collectively designed to assess various aspects of multimodal intelligence. For image understanding tasks, we conduct experiments on eight benchmarks: MME [Fu2023mme], ScienceQA [lu2022sqa], GQA [hudson2019gqa], POPE [li2023pope], MMBench and MMBench-CN [Liu2023mmbench], VizWiz [gurari2018vizwiz], and VQAv2 [goyal2017vqav2]. For video understanding tasks, we evaluate our method on four benchmarks: VideoMME [fu2024videomme], EgoSchema [mangalam2023egoschema], MLVU [zhou2024mlvu], and LongVideoBench [wu2024longvideobench].
@@ 1.1 Benchmarks
GQA is a benchmark designed to evaluate a model’s structured understanding and reasoning capabilities over visual scenes. It is built upon three key components: images, scene graphs, and carefully constructed questions. Each image is accompanied by a scene graph derived from the Visual Genome dataset [Krishna2016visualgc], which provides detailed object-level annotations, attributes, and relationships within the scene. The questions are automatically generated from these graphs to ensure semantic clarity and logical consistency, enabling fine-grained assessment of a model’s reasoning ability. Following standard practice, we report accuracy on the test-dev set, which contains 12,578 image-question pairs.
@@
POPE is a benchmark designed to assess object hallucination in vision-language models through binary questions about the presence of specific objects in images. The images are sourced from the MSCOCO dataset [Lin2014microsoftcc], and evaluation is based on the average F1 score across three sampling strategies, using a total of 8,910 image-question pairs.
@@
MMBench is a hierarchical benchmark designed to comprehensively evaluate multimodal model capabilities across three levels: perception and reasoning (L1), six sub-skills (L2), and 20 specific tasks (L3). Each task consists of multiple-choice questions. The benchmark is available in both English and Chinese, with 4,377 and 4,329 image-question pairs, respectively. We use both MMBench and MMBench-CN for evaluation.
@@
VQAv2 is a benchmark designed to evaluate a model’s visual recognition and reasoning capabilities through open-ended questions grounded in real-world images. It contains 265,016 images from the MSCOCO dataset [Lin2014microsoftcc], with each image paired with at least three questions. To mitigate bias, the dataset adopts an adversarially balanced design, ensuring that each question appears with multiple images leading to different answers. Each question is annotated with ten human-provided answers. We use the test-dev set for evaluation, which includes 107,394 image-question pairs, with scoring based on standard automatic metrics.
@@
EgoSchema is a diagnostic benchmark designed to evaluate the very long-form video-language understanding capabilities of LVLMs. Derived from the Ego4D dataset [Grauman2021ego4dat], it comprises over 5,000 human-curated multiple-choice question-answer pairs spanning more than 250 hours of egocentric video footage, covering a broad range of natural human activities and behaviors. Each question is based on a three-minute-long video clip and requires selecting the correct answer from five options. In our experiments, we evaluate on the EgoSchema test set, which contains 5,031 video-question pairs.
@@ 1.2 Backbone Models
LLaVA-NeXT [liu2024llavanext] is also referred to as LLaVA-1.6, extending LLaVA-1.5 [liu2024llava15] by introducing a dynamic high-resolution processing strategy that enhances performance in tasks requiring visual reasoning, OCR, and world knowledge. In contrast to the fixed resolution scaling used in LLaVA-1.5, LLaVA-NeXT adaptively adjusts the input resolution by selecting an optimal aspect ratio based on the original image. The resolution can be increased by up to 4×. Notably, this enhancement is achieved without modifying the visual encoder. Instead, each high-resolution image is divided into multiple sub-images of the original size, which are independently encoded and then concatenated before being passed to the language model. All experiments in this study are based on the 7B and 13B versions of LLaVA-NeXT.
Qwen2.5-VL.
Qwen2.5-VL [bai2025qwen25vl] is the flagship vision-language model in the Qwen family, featuring significant improvements in visual reasoning, localization, document understanding, and long-video comprehension. It supports object localization via bounding boxes or points and can output structured data (e.g. JSON) from documents, forms, tables, and diagrams. To handle complex visual inputs, Qwen-2.5-VL employs dynamic resolution processing and absolute time encoding, which allow it to process variable-resolution images and long-range videos without conventional resizing or normalization. A native dynamic-resolution ViT architecture with windowed attention is trained from scratch to balance efficiency and perceptual fidelity. In this work, we use the 7B instruction-tuned variant, Qwen-2.5-VL-7B-Instruct, for experiments.
InternVL2.
The InternVL2 [internvl2] series provides a family of multimodal large language models (MLLMs) available in multiple sizes (e.g., 1B–8B–76B+) and instruction-tuned variants. It is trained with long-context modeling to support not only single-image inputs but also multi-image and video comprehension. InternVL2 offers broad capability coverage, including document, chart, and OCR understanding, visual reasoning, grounding, and multi-image or video comprehension, while maintaining a consistent architecture across different model scales. In our experiments, we employ the 8B instruction-tuned model, InternVL2-8B.
1.3 Comparison Methods
@@
FastV [chen2024fastv] is a training-free method that reduces computational overhead in vision-language models by performing early-stage visual token pruning. It identifies and removes the least relevant tokens after the second LLM layer by averaging attention scores.
@@
SparseVLM [zhang2024sparsevlm] ranks the importance of both visual and textual tokens based on cross-modal attention, and introduces adaptive sparsity ratios along with a token recycling strategy to better utilize discarded information.
@@
FasterVLM [zhang2024fastervlm] leverages attention from the [CLS] token to visual tokens as an importance indicator, pruning the less relevant visual tokens accordingly.
@@
GlobalCom2 [liu2025globalcom2] is designed for high-resolution image understanding tasks that receive both a global thumbnail and multiple local crops. The thumbnail provides global contextual guidance to guide the compression of each crop in a task-specific manner.
DivPrune.
DivPrune [alvar2025divprune] formulates visual token retention as a min–max diversity problem, employing a greedy algorithm that iteratively selects tokens most dissimilar to those already chosen.
2 Theoretical Perspective: Orthogonality of Intra- and Cross-Modal Redundancy
@@
$\mathcal{V}=\operatorname{Span}(W_{V}),\hskip 28.80008pt\mathcal{T}=\operatorname{Span}(W_{T}),$
@@
$W_{V}^{\top}W_{T}=0\hskip 28.80008pt(\mathcal{V}\;\perp\;\mathcal{T}).$
@@
Proof of (12). Set $X_{ij}=\kappa(v_{i},v_{j})$ and $Y_{k}=\rho(v_{k},T)$. By Eqs.˜9 and 10,
$D_{\kappa}(V)=\frac{1}{N^{2}}\!\sum_{i\neq j}X_{ij},\hskip 28.80008ptR_{\rho}(V,T)=\frac{1}{N}\!\sum_{k=1}^{N}Y_{k}.$
@@
3 More Ablation Studies
3.1 Pivot Token Selection Strategies
Method MME ScienceQA GQA POPE Average
Original 1519.6 72.0 64.2 87.7 100%
Retention Ratio = 25%
[CLS] 1504.3 70.7 63.3 87.5 98.9%
Random 1500.0 70.6 63.2 87.1 98.6%
Center 1502.7 70.6 63.3 87.4 98.8%
Farthest 1500.3 70.7 63.2 87.4 98.7%
Retention Ratio = 10%
[CLS] 1469.3 70.5 59.4 85.0 96.0%
Random 1463.7 70.2 59.4 85.0 95.8%
Center 1455.7 70.1 59.5 84.8 95.6%
Farthest 1461.2 70.2 59.5 85.1 95.8%
Table 1: Ablations on Pivot Token Selection Strategy. All results are based on LLaVA-NeXT-7B. “[CLS]” = token with highest attention to encoder [CLS]; “Center” = token nearest to mean visual feature; “Farthest” = farthest token from mean; “Random” = random token.
We conduct an ablation study on different pivot token selection strategies used in the diversity-driven reduction stage. As shown in Table˜1, selecting the token with the highest attention to the encoder [CLS] token yields the best performance, while choosing the token nearest or farthest from the mean visual feature performs less effectively. Interestingly, randomly selecting a pivot token achieves comparable performance, suggesting that this strategy can serve as a practical alternative for MLLMs whose encoders do not contain a [CLS] token, thereby making ToDRE more transferable across different model architectures.
3.2 Threshold in Relevance-driven Visual Token Reduction
@@
[…]
0.03 79:22 1519.6 71.7 64.2 87.6 99.9%
0.05 73:24 1530.2 71.7 64.2 87.6 100.0%
0.10 (Ours) 72:35 1530.4 71.7 64.2 87.7 100.1%
0.15 72:25 1524.0 71.7 64.2 87.6 100.0%
Table 2: Ablation study on threshold $\bm{\tau}$ in relevance-driven visual token reduction. When both attention ratios $\alpha_{t\to v}$ and $\alpha_{v\to t}$ fall below the threshold $\tau$, all remaining visual tokens are removed from the LLM input.
We conduct an ablation study on the threshold $\tau$ used in the relevance-driven visual token reduction strategy, as shown in Table˜2. Varying $\tau$ controls the aggressiveness of token pruning based on the measured attention ratios $\alpha_{t\to v}$ and $\alpha_{v\to t}$. A larger $\tau$ leads to more extreme pruning but may sacrifice accuracy, while a smaller $\tau$ makes token pruning more conservative and thus improves performance at the cost of increased computational burden.
We choose $\tau=0.10$ as our default setting, which yields the optimal trade-off between efficiency and performance—maintaining 100.1% average accuracy while reducing inference time by over 4 minutes compared to the uncompressed baseline.
3.3 Layers in Relevance-driven Visual Token Reduction